# Can Watermarking Large Language Models Prevent Copyrighted Text Generation and Hide Training Data?

### v1

@@ 5.1 Empirical evaluation
In Table 4, we show the AUC for the unwatermarked and watermarked models using the UMD scheme, as well as the drop between the two. We observe that watermarking reduces the AUC (drop shown in bold in the table) by up to $14.2\%$ across $4$ detection methods and $5$ LLMs. All the experiments on watermarked models are run with $5$ different seeds and we report the mean and standard deviation of the results. We also report the AUC drop, which is computed by the difference between the AUC for the unwatermarked model and the mean AUC over the $5$ runs for the watermarked model. Additionally, while the experiments from Table 4 are conducted on WikiMIA-256, we observe similar trends for WikiMIA-32, WikiMIA-64, and WikiMIA-128 in Appendix A.2. We also study the impact of the watermark’s strength on the AUC drop for Llama-30B and GPT-Neo-20B in Figure 3 and for the other models in Figure 5 from the Appendix A.2. Note that we considered WikiMIA-256 for these experiments. We observe that higher watermark strengths generally induce larger AUC drops.
@@
Llama-30B Llama-13B
$85.4\%$ $68.2\%$
PPL $84.7\pm 1.4\%$ $67.6\pm 2.5\%$
0.7% 0.6%
$87.9\%$ $77.6\%$
Lowercase $80.9\pm 3.1\%$ $67.2\pm 4.0\%$
7.0% 10.4%
$82.5\%$ $62.5\%$
Zlib $77.8\pm 1.2\%$ $57.1\pm 2.0\%$
4.7% 5.4%
$85.1\%$ $70.2\%$
Min-KProb $85.0\pm 1.0\%$ $68.5\pm 0.1\%$
0.1% 1.7%
In addition to the $4$ detection methods, we also consider Smaller Ref attack, which we include in Table 12 of Appendix A.2. We consider different variations, including an unwatermarked reference model and a watermarked one with a similar strength but a different seed or with both strength and seed changed in comparison to the watermarked target model. The baseline is an unwatermarked model with an unwatermarked reference model. We observe the AUC drops in all scenarios (up to $16.4\%$), which is consistent with our previous findings.
We also experiment with several percentages of green tokens for a fixed watermark strength of $\delta=10$. We show the results in Table 13 of Appendix A.2. We observe that for all models, in at least $80\%$ of the cases all of the attacks’ AUCs are negatively affected (positive drop value), suggesting that, in general, finding a watermarking scheme that reduces the success rates of the current MIAs is not a difficult task. Note that the experiments are run on WikiMIA for UMD scheme and the results are averaged over $5$ watermark seeds.
@@
6 Improving Detection Performance with Adaptive Min-KProb
@@
This section demonstrates how an informed, adaptive attacker can improve the success rate of a recent MIA, Min-KProb. Our main idea is that an attacker with knowledge of the watermarking technique (including green-red token lists and watermark’s strength $\delta$) can readjust token probabilities. This is possible even without additional information about the logit distribution, relying solely on the probability of each token from the target sample given the preceding tokens. Our approach relies on two key assumptions. First, knowledge of the watermarking scheme, which aligns with assumptions made in prior work on public watermark detection [20]. Second, access to the probability of each token in a sample, given the previous tokens—an assumption also made by the Min-KProb method [31].
@@
Our method described in Algorithm 1 is based on the observation that if the denominator of softmax function (i.e., $\sum_{i}e^{z_{i}}$, where $z_{i}$ is the logit for the $i$-th vocabulary) does not vary significantly when generating samples with the watermarked model (and similarly for the unwatermarked model), then we can readjust the probabilities of the green tokens by “removing” the bias $\delta$. More precisely, assuming the approximation for the denominator of softmax is good, then the probability for each token $t_{i}$ in an unwatermarked model will be around $\frac{e^{L_{i}}}{c}$, where $L_{i}$ is the logit corresponding to the token $t_{i}$ and $c$ is a constant. However, for a watermarked model, if the token $t_{i}$ is green, then the probability would be approximated by $\frac{e^{L_{i}+\delta}}{d}$, where $d$ is again a constant, while in the case $t_{i}$ is red the probability will be around $\frac{e^{L_{i}}}{d}$. To compensate for the bias introduced by watermarking, we divide the probability of green tokens by $e^{\delta}$ and this way we end up with probabilities that are just a scaled (by $\frac{c}{d}$) version of the probabilities from the unwatermarked model. The scaling factor will not affect the orders between the samples when computing the average of the minimum Klog-probabilities as long as the tested sentences are approximately the same length, which is an assumption made by Shi et al. [31] as well.
@@
Despite the strong assumption we assumed regarding the approximation of the denominator, empirical results show that our method effectively improves the success rate of Min-Kunder watermarking. The results in Table 5, averaged over $5$ runs, demonstrate that our method improves over the baseline in $95\%$ of the cases, and the increase is as high as $4.8\%$.
@@
Takeaways. We demonstrate that an adaptive attacker can leverage the knowledge of a watermarking scheme to increase the success rate of a recent MIA, Min-KProb.
@@
WikiMIA Not adapt. $66.2\%$ $67.1\%$ $64.5\%$ $61.0\%$ $55.7\%$
32 Adapt. 68.5% 71.3% 66.3% $61.0\%$ 59.1%
WikiMIA Not adapt. $64.4\%$ $67.7\%$ $62.8\%$ $59.8\%$ $55.3\%$
64 Adapt. 67.3% 72.0% 64.9% 60.6% 57.4%
WikiMIA Not adapt. $70.0\%$ $73.0\%$ $68.9\%$ $64.8\%$ $59.2\%$
128 Adapt. 73.1% 75.9% 71.0% 66.4% 64.0%
WikiMIA Not adapt. $70.5\%$ $76.2\%$ $70.4\%$ $69.5\%$ $63.1\%$
256 Adapt. 71.3% 78.2% 72.4% $\textbf{70.7}\%$ 66.2%
7 Conclusion and Discussion
Watermarking LLMs has unintended consequences on methods towards copyright protection. Our experiments demonstrate that while watermarking may be a promising solution to prevent copyrighted text generation, watermarking also complicates membership inference attacks that may be employed to detect copyright abuses. Watermarking can be a double-edged sword for copyright regulators since it promotes compliance during generation time, while making training time copyright violations harder to detect. We hope our work further the discussion around watermarking and copyright issues for LLMs.
Limitations & Future Work. Our work considers only decoding time watermarking techniques, future work may benefit from studying other types of watermarking methods. Our proposed method for improving MIAs’ success rate on watermarked models makes strong assumptions on the watermarking scheme, which may not always be satisfied despite empirical improvements in our experiments. Our observations on the deterioration of MIAs’ success suggests that for copyright violation auditing, an unwatermarked model or the watermarking scheme may be needed. We encourage the community to further refine adaptive methods to ensure robust copyright protection and data privacy, and consider the interactions of different methods on downstream legal concerns.
Acknowledgements
Panaitescu-Liess, Che, An, Xu, Pathmanathan, Chakraborty, Zhu, and Huang are supported by DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) 80321, National Science Foundation NSF-IIS-2147276 FAI, DOD-ONR-Office of Naval Research under award number N00014-22-1-2335, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, DOD-DARPA-Defense Advanced Research Projects Agency Guaranteeing AI Robustness against Deception (GARD) HR00112020007, Adobe, Capital One and JP Morgan faculty fellowships.
This work was made possible by the ONR MURI program and the AFOSR MURI program. Commercial support was provided by Capital One Bank, the Amazon Research Award program, and Open Philanthropy. Further support was provided by the National Science Foundation (IIS-2212182), and by the NSF TRAILS Institute (2229885).
Appendix A Appendix
A.1 Additional experiments on verbatim memorization on WikiMIA
Table 6: Measuring the reduction in verbatim memorization of copyrighted texts on WikiMIA-64. We report the relative increase in perplexity between the watermarked and unwatermarked model, so larger values correspond to less memorization. Note that "P." stands for "prompt length".
@@ A.1 Additional experiments on verbatim memorization on WikiMIA
Table 7: Measuring the reduction in verbatim memorization of copyrighted texts on WikiMIA-128. We report the relative increase in perplexity between the watermarked and unwatermarked model, so larger values correspond to less memorization. Note that "P." stands for "prompt length".
@@
Table 8: Measuring the reduction in verbatim memorization of copyrighted texts on WikiMIA-256. We report the relative increase in perplexity between the watermarked and unwatermarked model, so larger values correspond to less memorization. Note that "P." stands for "prompt length".
@@
A.2 Additional experiments on pretraining data detection on WikiMIA
@@
A.3 Additional experiments on BookMIA
In this section, we conduct experiments using models finetuned on a subset of BookMIA, which we refer to as BookMIA-2. To build BookMIA-2, we first select only the samples that were not part of the training set of any model that we consider (labeled as 0 by [31]). Then, we randomly select half of them as finetuning data (referred to as seen samples) and keep the other half as unseen samples. Note that for BookMIA-2, there would not be a distribution difference between the seen and unseen samples [12]. We also consider duplicating a sample from the training set of BookMIA-2 to have more fine-grained control over the memorization of that sample. We use Llama-7B in all the experiments from this section.
@@ A.3 Additional experiments on BookMIA
Approximate Memorization. Informally, we consider a training sample approximately memorized by a model if, given its prefix, it is possible to generate a completion that is similar enough to the ground truth completion. In our experiments, we consider Normalized Edit Similarity (referred to as edit similarity from now on) and BLEU score as similarity measures, as in [17]. Note that we consider both word-level and token-level variants for the BLEU score. The range for each metric is between 0 and 1, where values close to 1 represent similar texts. In all experiments, since all the samples are 512 words long, we consider the first 256 words as the prefix and the last 256 words as the ground truth completion. We show the results in Figure 6 averaged over 20 runs with different seeds. Note that the duplication factor (shown on x-axis) represents the number of times the target copyrighted text is duplicated. We observe that for high levels of memorization, a strong watermark significantly reduces the similarity between the generated completion and the ground truth (copyrighted) one.
@@
MIA. We also study the effect of the watermark on the effectiveness of MIAs for copyrighted training data detection (on BookMIA-2, without duplicated samples). We show the results in Table 15 and observe that watermarking negatively affects MIAs’ success rate, which is consistent with our previous findings (from Section 5). Finally, we also run our adaptive method from Section 6 and observe an improvement of $0.9\%$ over Min-KProb.
@@
A.4 Proofs for the theoretical results
In this section we present a complete proof for Theorem 2 from Section 4.2.
Theorem 2 (Restated) Let $f$ be a LLM and $f_{W}$ its watermarked version with a “soft” UMD scheme and let $\epsilon\in(0,\frac{1}{4})$. Let $s=t_{1}\oplus t_{2}\oplus...\oplus t_{n}\in C_{D}$ be a copyrighted sample. We consider $\gamma=0.5$. We denote the output of the softmax layer of $f$ for generating the token $t_{i}$ as $\frac{a_{i}}{d_{i}+a_{i}}$ and in the case of $f_{W}$, we denote it by $\frac{a_{i}\cdot e^{\delta}}{b_{i}^{{}^{\prime}}+c_{i}^{{}^{\prime}}\cdot e^{\delta}+a_{i}\cdot e^{\delta}}$ (if $t_{i}$ is on the green list) and $\frac{a_{i}}{b_{i}^{{}^{\prime\prime}}+c_{i}^{{}^{\prime\prime}}\cdot e^{\delta}+a_{i}}$ (if $t_{i}$ is on the red list), where $a_{i}$ is the exponential of the logit value corresponding to the token $t_{i}$ and $b_{i}^{{}^{\prime}},b_{i}^{{}^{\prime\prime}}$ and $c_{i}^{{}^{\prime}},c_{i}^{{}^{\prime\prime}}$ are the sum of the exponentials of the logits corresponding to other tokens that are on the red list and green list, respectively. We assume that $\frac{x}{a_{i}}<M=\frac{1-4\epsilon}{1+4\epsilon}$, for all $x\in\{d_{i},b_{i}^{{}^{\prime}},b_{i}^{{}^{\prime\prime}},c_{i}^{{}^{\prime}},c_{i}^{{}^{\prime\prime}}\}$ which would restrict $f$ to be relatively confident in its predictions for each token $t_{i}$. Then, we can always find a $\delta$ (strength) for the watermarking scheme such that the probability of generating $s$ is reduced by at least $(1+\frac{2\epsilon}{2\epsilon+1})^{n}$ times in comparison to the case of the unwatermarked model.

### v3

@@ Introduction
In Table 4, we show the AUC for the unwatermarked and watermarked models using the UMD scheme, as well as the drop between the two. We observe that watermarking reduces the AUC (drop shown in bold in the table) by up to $14.2\%$ across $4$ detection methods and $5$ LLMs. All the experiments on watermarked models are run with $5$ different seeds and we report the mean and standard deviation of the results. We also report the AUC drop, which is computed by the difference between the AUC for the unwatermarked model and the mean AUC over the $5$ runs for the watermarked model. Additionally, while the experiments from Table 4 are conducted on WikiMIA-256, we observe similar trends for WikiMIA-32, WikiMIA-64, and WikiMIA-128 in the appendix. We also study the impact of the watermark’s strength on the AUC drop for Llama-30B in Figure 4 and for the other models in Figure 6 from the appendix. Note that we considered WikiMIA-256 for these experiments. We observe that higher watermark strengths generally induce larger AUC drops.
Llama-30B Llama-13B
PPL $85.4\%$ $68.2\%$
$84.7\pm 1.4\%$ $67.6\pm 2.5\%$
0.7% 0.6%
Lowercase $87.9\%$ $77.6\%$
$80.9\pm 3.1\%$ $67.2\pm 4.0\%$
7.0% 10.4%
Zlib $82.5\%$ $62.5\%$
$77.8\pm 1.2\%$ $57.1\pm 2.0\%$
4.7% 5.4%
Min-KProb $85.1\%$ $70.2\%$
$85.0\pm 1.0\%$ $68.5\pm 0.1\%$
0.1% 1.7%
@@
In addition to the $4$ detection methods, we also consider Smaller Ref attack, which we include in Table 13 of the appendix. We consider different variations, including an unwatermarked reference model and a watermarked one with a similar strength but a different seed or with both strength and seed changed in comparison to the watermarked target model. The baseline is an unwatermarked model with an unwatermarked reference model. We observe the AUC drops in all scenarios (up to $16.4\%$), which is consistent with our previous findings.
We also experiment with several percentages of green tokens for a fixed watermark strength of $\delta=10$. We show the results in Table 14 of the appendix. We observe that for all models, in at least $80\%$ of the cases all of the attacks’ AUCs are negatively affected (positive drop value), suggesting that, in general, finding a watermarking scheme that reduces the success rates of the current MIAs is not a difficult task. Note that the experiments are run on WikiMIA for UMD scheme and the results are averaged over $5$ watermark seeds.
Improving Detection Performance with Adaptive Min-KProb
@@
This section demonstrates how an informed, adaptive attacker can improve the success rate of a recent MIA, Min-KProb. Our main idea is that an attacker with knowledge of the watermarking technique (including green-red token lists and watermark’s strength $\delta$) can readjust token probabilities. This is possible even without additional information about the logit distribution, relying solely on the probability of each token from the target sample given the preceding tokens. Our approach relies on two key assumptions. First, knowledge of the watermarking scheme, which aligns with assumptions made in prior work on public watermark detection (Kirchenbauer et al. 2023). Second, access to the probability of each token in a sample, given the previous tokens—an assumption also made by the Min-KProb method (Shi et al. 2023).
Threat model. (1) The attacker’s goal is to infer whether specific samples are part of the training set or not. In our setting, the attacker is not malicious, as the goal is to detect copyright violations. (2) Regarding the attacker’s knowledge, we assume the attacker knows the watermarking method and its parameters (green and red lists, and the watermark strength), which aligns with the assumption made in prior work by Kirchenbauer et al. (2023) for public watermark detection. (3) As for the attacker’s capabilities, we assume they can access the probabilities for each token in the given samples, similar to what a copyright auditor may have access to. This also mirrors the assumption made by Shi et al. (2023) in the context of training data detection.
@@
Our method described in Algorithm 1 is based on the observation that if the denominator of softmax function (i.e., $\sum_{i}e^{z_{i}}$, where $z_{i}$ is the logit for the $i$-th vocabulary) does not vary significantly when generating samples with the watermarked model (and similarly for the unwatermarked model), then we can readjust the probabilities of the green tokens by “removing” the bias $\delta$. More precisely, assuming the approximation for the denominator of softmax is good, then the probability for each token $t_{i}$ in an unwatermarked model will be around $\frac{e^{L_{i}}}{c}$, where $L_{i}$ is the logit corresponding to the token $t_{i}$ and $c$ is a constant. However, for a watermarked model, if the token $t_{i}$ is green, then the probability would be approximated by $\frac{e^{L_{i}+\delta}}{d}$, where $d$ is again a constant, while in the case $t_{i}$ is red the probability will be around $\frac{e^{L_{i}}}{d}$. To compensate for the bias introduced by watermarking, we divide the probability of green tokens by $e^{\delta}$ and this way we end up with probabilities that are just a scaled (by $\frac{c}{d}$) version of the probabilities from the unwatermarked model. The scaling factor will not affect the orders between the samples when computing the average of the minimum Klog-probabilities as long as the tested sentences are approximately the same length, which is an assumption made by Shi et al. (2023) as well.
@@
Despite the strong assumption we assumed regarding the approximation of the denominator, empirical results show that our method effectively improves the success rate of Min-Kunder watermarking. We show results for two LLMs in Table 5 and include the complete results for 5 LLMs in Table 17 from the appendix. We observe that our method improves over the baseline in $95\%$ of the cases, and the increase is as high as $4.8\%$ (averaged over 5 runs).
Finally, we also consider adaptive versions of the Lowercase and Zlib methods. Our findings show that these adaptive methods outperform the baselines in at least 80% of cases. Detailed results are provided in the appendix.
@@
Takeaways. We demonstrate that an adaptive attacker can leverage the knowledge of a watermarking scheme to increase the success rates of recent MIAs.
@@
Llama-30B Llama-13B
WikiMIA Not adapt. $66.2\%$ $64.5\%$
32 Adapt. 68.5% 66.3%
WikiMIA Not adapt. $64.4\%$ $62.8\%$
64 Adapt. 67.3% 64.9%
WikiMIA Not adapt. $70.0\%$ $68.9\%$
128 Adapt. 73.1% 71.0%
WikiMIA Not adapt. $70.5\%$ $70.4\%$
256 Adapt. 71.3% 72.4%
@@
Conclusion and Discussion
Watermarking LLMs has unintended consequences on methods towards copyright protection. Our experiments demonstrate that while watermarking may be a promising solution to prevent copyrighted text generation, watermarking also complicates membership inference attacks that may be employed to detect copyright abuses. Watermarking can be a double-edged sword for copyright regulators since it promotes compliance during generation time, while making training time copyright violations harder to detect. We hope our work furthers the discussion around watermarking and copyright issues for LLMs.
Acknowledgements
Panaitescu-Liess, Che, An, Xu, Pathmanathan, Chakraborty, Zhu, and Huang are supported by DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) 80321, National Science Foundation NSF-IIS-2147276 FAI, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, Adobe, Capital One and JP Morgan faculty fellowships.
Appendix A Appendix
Additional experiments on verbatim memorization on WikiMIA
@@ Appendix A Appendix
P. Min. Avg. Min. Avg. Min. Avg. Min. Avg. Min. Avg.
0 $3.3$ $31.2$ $3.7$ $52.1$ $4.9$ $34.3$ $11.4$ $61.3$ $10.4$ $64.5$
UMD 10 $2.8$ $28.7$ $2.2$ $52.1$ $3.5$ $31.9$ $8.8$ $63.7$ $8.3$ $67.7$
20 $2.4$ $30.1$ $1.8$ $66.0$ $3.5$ $33.4$ $5.0$ $74.0$ $7.0$ $84.4$
0 $4.1$ $34.1$ $4.4$ $54.1$ $5.0$ $36.6$ $14.3$ $74.5$ $11.5$ $66.1$
Unigram 10 $3.0$ $31.7$ $2.8$ $52.5$ $4.0$ $34.3$ $11.8$ $73.6$ $9.8$ $70.2$
20 $2.4$ $31.5$ $2.0$ $56.4$ $3.4$ $34.0$ $6.6$ $79.1$ $5.8$ $81.4$
0 $4.0$ $34.3$ $4.9$ $51.1$ $5.5$ $34.7$ $8.0$ $62.3$ $7.4$ $60.6$
Random 10 $2.6$ $31.4$ $3.1$ $51.4$ $3.6$ $31.7$ $5.5$ $62.9$ $6.3$ $64.0$
20 $2.1$ $31.8$ $1.2$ $59.3$ $2.8$ $32.9$ $4.7$ $78.6$ $3.9$ $73.7$
Table 7: Measuring the reduction in verbatim memorization of training texts on WikiMIA-64. We report the relative increase in both the minimum and average perplexity between the watermarked and unwatermarked models, where larger values correspond to less memorization. Note that “P.” stands for “prompt length”.
@@
Table 8: Measuring the reduction in verbatim memorization of training texts on WikiMIA-128. We report the relative increase in both the minimum and average perplexity between the watermarked and unwatermarked models, where larger values correspond to less memorization. Note that “P.” stands for “prompt length”.
@@
Table 9: Measuring the reduction in verbatim memorization of training texts on WikiMIA-256. We report the relative increase in both the minimum and average perplexity between the watermarked and unwatermarked models, where larger values correspond to less memorization. Note that “P.” stands for “prompt length”.
@@
Additional experiments on pretraining data detection on WikiMIA
@@
Additional experiments on BookMIA
In this section, we conduct experiments using models finetuned on a subset of BookMIA, which we refer to as BookMIA-2. To build BookMIA-2, we first select only the samples that were not part of the training set of any model that we consider (labeled as 0 by (Shi et al. 2023)). Then, we randomly select half of them as finetuning data (referred to as seen samples) and keep the other half as unseen samples. Note that for BookMIA-2, there would not be a distribution difference between the seen and unseen samples (Das, Zhang, and Tramèr 2024). We also consider duplicating a sample from the training set of BookMIA-2 to have more fine-grained control over the memorization of that sample. We use Llama-7B in all the experiments from this section.
@@
MIA. We also study the effect of the watermark on the effectiveness of MIAs for copyrighted training data detection (on BookMIA-2, without duplicated samples). We show the results in Table 16 and observe that watermarking negatively affects MIAs’ success rate, which is consistent with our previous findings (from Section Impact of Watermarking on Pretraining Data Detection). Finally, we also run our adaptive method from Section Improving Detection Performance with Adaptive Min-KProb and observe an improvement of $0.9\%$ over Min-KProb.
@@
Theoretical analysis
Theorem 1
For an LLM watermarked using a “hard” UMD scheme with a percentage of $\gamma$ green tokens, then the probability of generating a copyrighted text is lower than $m\cdot\gamma^{n}$.
Proof. Given one sample $s=t_{1}\oplus t_{2}\oplus...\oplus t_{n}\in C_{D}$. For a “hard” watermarking scheme, the probability $P(s)$ of generating $s$ is smaller than the probability of each token $t_{i}$ to be on a green list. So, $P(s)<\gamma^{n}$. The probability of not generating any $s_{j}\in C_{D}$ is $P(\neg s_{1}\wedge\neg s_{2}\wedge...\wedge\neg s_{m})=\prod_{i=\overline{1,m}}(1-P(s_{i}))>(1-\gamma^{n})^{m}>1-m\gamma^{n}.$ Note that we used Bernoulli’s inequality at the end. So, the probability of generating at least one copyrighted text is lower than $1-(1-m\cdot\gamma^{n})$ and hence lower than $m\cdot\gamma^{n}$.
Example. Let’s consider a “hard” UMD watermarking scheme with $\gamma=0.5$. Let’s assume each copyrighted text is $100$ tokens, the model was trained on a dataset containing $10^{9}$ copyrighted texts. The probability to generate a copyrighted text is $<10^{9}\cdot 0.5^{100}=\frac{10^{9}}{2^{100}}=\frac{1000^{3}}{1024^{10}}<\frac{1000^{3}}{1000^{10}}=1000^{-7}=10^{-21}$ and hence very low.
Theorem 2
@@
Observation. Since the probability is reduced by at least $(1+\frac{2\epsilon}{2\epsilon+1})^{n}$ times in Theorem 2 then the probability of generating $s$ is lower than $(\frac{2\epsilon+1}{4\epsilon+1})^{n}$ (as the maximum probability of generating with the unwatermarked model is $1$). Hence, as in Theorem 1, we observe that the probability of generating a copyrighted text is lower than $m\cdot(\frac{2\epsilon+1}{4\epsilon+1})^{n}$.
Takeaways. Our theoretical analysis demonstrates that watermarking significantly reduces the probability of generating copyrighted text verbatim. For both a “hard” and “soft” UMD scheme, the upper bound for the likelihood of producing copyrighted content decreases exponentially with the length of the copyrighted texts.
Metrics used in MIAs
In this section we provide short formulas for several metrics used by MIAs such as Smaller Ref, Lowercase and Zlib (Carlini et al. 2021).
Let $f$ be the model the MIA is applied to, $g$ be a smaller model, $x$ be a sample (string), log be the natural logarithm function, perplexity_h(x) be a function that computes the perplexity of a model $h$ on $x$, lowercase be a function that maps a string to its lowercased version, and zlib(x) be a function that computes the zlib entropy of a string $x$.
Smaller Ref uses the metric $\frac{\texttt{log({perplexity\_f(x)})}}{\texttt{log({perplexity\_g(x)})}}$.
Lowercase uses the metric $\frac{\texttt{log({perplexity\_f(x)})}}{\texttt{log({perplexity\_f({lowercase}(x))})}}$.
Zlib uses the metric $\frac{\texttt{log({perplexity\_f(x)})}}{\texttt{zlib}(x)}$.
Adaptive Min-KProb
Llama-30B NeoX-20B Llama-13B Pythia-2.8B OPT-2.7B
WikiMIA Not adapt. $66.2\%$ $67.1\%$ $64.5\%$ $61.0\%$ $55.7\%$
32 Adapt. 68.5% 71.3% 66.3% $61.0\%$ 59.1%
WikiMIA Not adapt. $64.4\%$ $67.7\%$ $62.8\%$ $59.8\%$ $55.3\%$
64 Adapt. 67.3% 72.0% 64.9% 60.6% 57.4%
WikiMIA Not adapt. $70.0\%$ $73.0\%$ $68.9\%$ $64.8\%$ $59.2\%$
128 Adapt. 73.1% 75.9% 71.0% 66.4% 64.0%
WikiMIA Not adapt. $70.5\%$ $76.2\%$ $70.4\%$ $69.5\%$ $63.1\%$
256 Adapt. 71.3% 78.2% 72.4% $\textbf{70.7}\%$ 66.2%
Adaptive Zlib and Adaptive Lowercase
We have also adapted Zlib, as well as Lowercase. In these adaptations, we use a method similar to Adaptive Min-KProb to approximate pre-watermarked probabilities, then we apply Zlib / Lowercase. The results are presented in Tables 18 and 19, where each double-cell shows the AUC score for the non-adaptive attack on top and the adaptive version on the bottom. We observe that adaptation improves the baseline in over 80% of cases (for Zlib) and over 90% of cases (for Lowercase). However, these adaptive methods appear slightly less effective compared to Min-KProb. We hypothesize that this may be due to them using all token probabilities rather than the minimum Ksubset, as in Min-KProb, which may increase error accumulation due to our token probability approximation.
Table 18: We show the AUC of Zlib (referred as “Not adapt.”) and our method (referred as “Adapt.”) when using UMD watermarking scheme. We highlight the cases when our method improves over the baseline.
Llama-30B NeoX-20B Llama-13B
WikiMIA Not adapt. 64.4% 63.2% 62.3%
32 Adapt. 66.8% 64.2% 64.9%
WikiMIA Not adapt. 62.4% 62.0% 59.9%
64 Adapt. 66.8% 63.6% 64.8%
WikiMIA Not adapt. 68.6% 66.3% 65.8%
128 Adapt. 71.4% 67.3% 69.1%
WikiMIA Not adapt. 72.0% 66.6% 71.6%
256 Adapt. 71.3% 67.1% 71.1%
Table 19: We show the AUC of Lowercase (referred as “Not adapt.”) and our method (referred as “Adapt.”) when using UMD watermarking scheme. We highlight the cases when our method improves over the baseline.
Llama-30B NeoX-20B Llama-13B
WikiMIA Not adapt. 54.9% 59.4% 54.2%
32 Adapt. 57.4% 62.0% 56.9%
WikiMIA Not adapt. 54.8% 56.8% 53.8%
64 Adapt. 56.7% 59.4% 56.0%
WikiMIA Not adapt. 55.9% 58.2% 55.1%
128 Adapt. 57.2% 59.4% 56.0%
WikiMIA Not adapt. 63.8% 55.4% 61.6%
256 Adapt. 63.5% 58.1% 62.3%
Examples of generated samples
Below we provide text samples generated from the unwatermarked model and from models with various UMD watermark strengths. All examples are generated from an empty string (free generation). These examples have relative perplexity increases that match or even exceed the corresponding averages in Figure 2.
w/o watermark: “Nonviolent communication, also called compassionate communication, is a way of relating to others based on values of human dignity, honesty, and empathy …”
w/ watermark (strength = 2): “The Latifa Hospital was established in Dubai, United Arab Emirates in March 1983 to provide advanced health care within Dubai …”
w/ watermark (strength = 4): “We all make New Year’s Resolutions and all of them start with “getting fitter” :) If it happens to be in your list, let me know …”
w/ watermark (strength = 6): “I’ll be the first to say that MTS doesn’t always achieve what it sets out to do at first look but you can see the software is reaching its potential …”
w/ watermark (strength = 8): “United Nations : Member states of UNO climate change conference agreed on Thursday that progress had to accelerate in implementing …”
w/ watermark (strength = 10): “Like our Facebook page to see when the weekend is set and to find out about local motorcycling news, events and …”
w/ watermark (strength = 12): “Your child will learn time management as well: it will learn how much time to spend doing each assignment and studying for tests to ensure they get them done …”
w/ watermark (strength = 14): “Grant is a shy four to six year old beagle boy with one blind eye and lacking much vision in both eyes. His family had been unable to …”
w/ watermark (strength = 16): “We are leading online doctor shopping solution company in Noida. It offers the best web solutions with goal orientated expertise in medical sector website development and marketing …”
Extended Related Work
Memorization. One cause of copyright issues is that machine learning models may memorize training data. Prior studies have observed that LLMs can memorize private information in training data, such as phone numbers and addresses (Karamolegkou et al. 2023; Carlini et al. 2019, 2021; Lee et al. 2021), leading to significant privacy and security concerns. To measure memorization, Carlini et al. (2021) proposes eidetic memorization, defining a string as memorized if it was present in the training data and it can be reproduced by a prompt. This definition, along with variations like exact and perfect memorization, has been widely adopted in subsequent studies (Tirumala et al. 2022; Kandpal, Wallace, and Raffel 2022). Carlini et al. (2022b) quantitatively measures memorization in LLMs as the fraction of extractable training data and finds that memorization significantly grows as model size scales and training examples are duplicated. To minimize memorization, Lee et al. (2021) and Kandpal, Wallace, and Raffel (2022) propose deduplicating training data, which also improves accuracy. Hans et al. (2024) proposes the Goldfish Loss as a training-time defense against verbatim memorization. Ippolito et al. (2023) proposes an inference time defense that perfectly prevents all verbatim memorization. However, it cannot prevent the leakage of training data due to the existence of many “style-transfer” prompts, suggesting it is a challenging open problem. Unlike the methods that we are studying in this paper, Ippolito et al. (2023) requires access to a complete set of copyrighted texts that the model was trained on. Memorization in the image domain has also been studied from various angles (Somepalli et al. 2023a, b; Carlini et al. 2023; Wen et al. 2024).
Membership Inference. As a proxy for measuring memorization, membership inference attacks (MIAs) predict whether or not a particular example was used to train the model (Shokri et al. 2017; Yeom et al. 2018; Bentley et al. 2020). Most membership inference attacks rely only on the model’s loss since the model is more likely to overfit an example if it is in the training data (Sablayrolles et al. 2019). Carlini et al. (2022a) trains shadow models to predict whether an example is from the training data. In the NLP domain, many works have focused on masked language models (Mireshghallah et al. 2022) and fine-tuning data detection (Song and Shmatikov 2019; Shejwalkar et al. 2021). Recently, Shi et al. (2023) studies pretraining data inference and introduced a detection method based on the hypothesis that unseen examples are likely to contain outlier words with low probabilities under the LLM. Zhang et al. (2024) approaches pretraining data detection by measuring how sharply peaked the likelihood is around the inputs. Duarte et al. (2024) proposes detecting copyrighted content in training data by probing the LLM with multiple-choice questions, whose options include both verbatim text and their paraphrases. Other methods include testing perplexity differences (Mattern et al. 2023) and providing provable guarantees of test set contamination without access to pretraining data or model weights (Oren et al. 2023).
Computing Infrastructure
All of our experiments were run on either three Nvidia RTX A6000 or four Nvidia RTX A5000 GPUs, using 128 GB of memory. We used the Transformers library (version 4.35.2) and PyTorch (version 2.1.0).
Limitations & Discussion
We emphasize that our claim is not that any watermarking method inherently prevents the generation of verbatim copyrighted text. Instead, we demonstrate that popular watermarking methods can unintentionally produce this effect. We believe that a distortion-free watermark, such as the one proposed by Kuditipudi et al. (2023), would not have this unintended effect. Our proposed method for improving MIAs’ success rate on watermarked models makes strong assumptions on the watermarking scheme, which may not always be satisfied despite empirical improvements in our experiments. Our observations on the deterioration of MIAs’ success suggests that for copyright violation auditing, an unwatermarked model or the watermarking scheme may be needed. We encourage the community to further refine adaptive methods to ensure robust copyright protection and data privacy, and consider the interactions of different methods on downstream legal concerns.